Suppose there are 5 red chips and 8 blue chips in a container. Two chips are selected out of the container without replacement. Find the probability of selecting a red chip on the second selection given that a blue was selected first
Answer by mathmate(429)


Question:
Suppose there are 5 red chips and 8 blue chips in a container. Two chips are selected out of the container without replacement. Find the probability of selecting a red chip on the second selection given that a blue was selected first

Solution:
Since one blue chip has already been selected (without replacement), there are 5 red and 7 blue left.
So the probability of selecting a red out of a total of 12 is 5/12.
Alternatively,
P(R|B)=P(R∩B)/P(B)=(5/13*8/12)/(8/13)=5/12 as before.